Suppose a recurrence relation


an=2an−1−an−2

where a1=7 and a2=10


can be represented in explicit formula, either as:

Formula 1:

an=pxn+qnxn

              or  

Formula 2:

an=pxn+qyn

 

where 

x

and

y

are roots of the characteristic equation.


Determine p and q


 Answer:

p :

q :


1
Expert's answer
2021-11-30T11:21:16-0500

Let us solve the characteristic equation of the recurrence relation "a_n=2a_{n\u22121}\u2212a_{n\u22122}," which is equivalent to "a_n-2a_{n\u22121}+a_{n\u22122}=0." It follows that the characteristic equation "x^2-2x+1=0" is equivalent to "(x-1)^2=0," and hence has the roots "x_1=x_2=1." It follows that the solution of the recurrence equation is "a_n=p\cdot 1^n+q\cdot n1^n=p+q\cdot n." Since "a_1=7" and "a_2=10," we conclude that "7=a_1=p+q" and "10=a_2=p+2q." Therefore, "q=3" and "p=4."

Answer: "p=4,\ q=3."